Project

General

Profile

Meeting 18 Jan 2013

Added by Tunnell, Christopher over 6 years ago

Agenda

  • Minutes and actions
  • Project manager report (Nichols)
  • MICE CM35: meeting goals etc.
  • preparations towards MICE Project Board
  • MICE data challenge (Colling)
  • Speakers board (Palladino)
  • future MICE CM
  • aob

Comments

Added by Tunnell, Christopher over 6 years ago

Email from Andy Nichols:

Alain,

Our Group is being sent home because of the weather and as I have to
cycle across country I might be some time! But I’ll phone from home if I
can.

Random PM comments for agenda item two:

·Firstly, great news about FC #1 - very gratifying to see so many people
persistently working through the problems.

·Roy Preece will has assumed the UK Project Manager role; this leaves a
vacancy again for the magnet systems level II Manager in the WBS

·Plans to develop and staff the operational and installation side of
MICE are being incorporated in FY13/14 UK plan

·Agreement reached with ISIS over minimal set of changes to the LH2
system for transition between system test and Step IV configuration.

·The above meeting exposed a number of worrying open engineering issues
with the liquid absorber and window integration, such as the sealing,
mounting and testing methods. This will be an agenda item for the CM35
engineering session; Wing Lau has been asked to report – I suspect that
this will identify some more work for us to do. Shigeru Ishimoto has
agreed to attend when he is required, which is good news.

·If the CERN magnet mapping team is required at Wang for scientific
reasons
, that’s fine, but we have to accept that the AFC is done first
and then the SS strictly in sequence and if that implies a schedule
delay, then we have to be realistic. We must also respect the priorities
of the CERN mapping group.

·An offer has been made to a suitable full time RF engineer for MICE at
DL, who will work initially with Andy Moss before coming to RAL to build
and lead an RF team there (assuming that he accepts of course)

·The global field mitigation baseline is established; physical work on
the West wall compressor is about to start and a detailed costed plan of
the rack/control room configuration is underway. Craig Macwaters has
been invited to the CM35 schedule discussion to report the field
mitigation milestones and to discuss how the project can be described in
the WBS

·My presumption is that the Step IV schedule baseline will be unchanged
from that at CM34.

Cheers,

Andy

Added by Tunnell, Christopher about 6 years ago

Present:

  • Maurizzio Bonesini
  • Chris Tunnell
  • Alain Blondel
  • Dan Kaplan
  • Alan Bross
  • John Cobb
  • Linda Coney

Apologies:

  • Andy Nichols

Review old minutes

  • Chris Rogers taken over organizing CM
  • Magnetic field mitigation for an afternoon
  • Blondel sent something to Courthold about field mitigation elements. CLOSED.
  • Forces between magnets in new configurations
  • Infinite combinations of magnetic currents with yokes
  • Which is worst? Unclear.
  • How much effort to invest?
  • Holger Witte is going to CM. For credible discussion of partial return yoke, can't disrupt Holger.
  • Mike Courthold only works 3 days a week

action: Procedures for magnetic measurements

  • John Cobb looked at Alain's work from previous meeting. Seemed happy with it.
  • How many measurements in each dimensions? What order to take measurements? Depends on how long it takes to move.
    ACTION: Blondel to send note to Victoria Blackmore, Cobb, and Felix Bergma with guesses on what points should be measured. Victoria is doing the fitting.

Who does the instrumentation of the absorber?

Define interface between Shigiru and RAL expert.

Project manager report

The following was emailed from Andy Nichols.

Firstly, great news about FC #1 - very gratifying to see so many people

persistently working through the problems.

Roy Preece will has assumed the UK Project Manager role; this leaves a vacancy again for the magnet systems level II Manager in the WBS

Need new L2 manager

Plans to develop and staff the operational and installation side of

MICE are being incorporated in FY13/14 UK plan

Agreement reached with ISIS over minimal set of changes to the LH2

system for transition between system test and Step IV configuration.

The above meeting exposed a number of worrying open engineering issues
with the liquid absorber and window integration, such as the sealing,
mounting and testing methods. This will be an agenda item for the CM35
engineering session; Wing Lau has been asked to report – I suspect that
this will identify some more work for us to do. Shigeru Ishimoto has
agreed to attend when he is required, which is good news.

If the CERN magnet mapping team is required at Wang for scientific
reasons, that’s fine, but we have to accept that the AFC is done first
and then the SS strictly in sequence and if that implies a schedule
delay, then we have to be realistic. We must also respect the priorities
of the CERN mapping group.

  • Adds schedule difficulty if we do AFC@RAL, send to Wang for SS, then back to RAL.
  • Bross: do we need to measure in hall? Need to demonstrate need more.
  • Measurement at Wang using MICE measurements system came from US SS review
  • Once mapped magnet well enough in one environment, should be able to extrapolate. If you don't believe this premise, then you have to measure in many places.
  • Could possibly train FNAL TD in how to use CERN system? Free up schedule constraints on CERN side?

An offer has been made to a suitable full time RF engineer for MICE at
DL, who will work initially with Andy Moss before coming to RAL to build
and lead an RF team there (assuming that he accepts of course)

The global field mitigation baseline is established; physical work on
the West wall compressor is about to start and a detailed costed plan of
the rack/control room configuration is underway. Craig Macwaters has
been invited to the CM35 schedule discussion to report the field
mitigation milestones and to discuss how the project can be described in
the WBS

My presumption is that the Step IV schedule baseline will be unchanged
from that at CM34.

Where are we with SS

  • Fully assembled
  • vacuum
  • Control issues mostly addresses (rest wait until DL comes out)
  • Cooldown Monday: Jan 18th
  • Cold by end of that week
  • Marginally questionable HTS leads replaced
  • So far, everything looks good.

MICE Data challenge

Nice idea.

Speaker's board

  • Three confirmed talks at IPAC13 Shanghai
  • APS April. Heidt? Lane?
  • Presented Mark and Victoria's work numerous times
  • Summer conferences.
  • Step 4 physics goals?

Future MICE CM

  • If RAL, need to plan ahead.